# From Concept to Reality: 5G Positioning with Open-Source Implementation of UL-TDoA in OpenAirInterface

### v1

@@ I Introduction
There are several examples on the use of OAI for positioning, in addition to HOP-5G testbed. In [3], OAI is combined with a multi-channel SDR and a custom 2x2 patch antenna array for enhanced cell-ID (E-CID) positioning in indoor scenarios, by using RTT and AoA estimates based on the demodulation reference signal (DM-RS). Thanks to the OAI PRS developments in the HOP-5G project [4], a testbed with three OAI base stations or gNBs and the OAI user equipment is deployed in [5] to exploit the PRS transmissions for DL-TDoA positioning. More recently, a novel framework is presented in [6] to estimate the RTT between a UE and a base station (gNB) with existing 5G NR transmissions.
@@ III-A UL-TDoA Positioning Procedure in OAI
The UL-TDoA positioning procedure begins with an external API generating a location information request for a target UE for the LMF. As shown in Figure  5 to initiate the location procedure, the external API (which can be an advanced API or a simple one-line command) sends the HTTP post request to determine-location API of LMF, where the request contains a data structure of type InputData [9, Section 6.1.6.2.2]. To enable the exchange of this message, we first developed the 3gpp-compliant LMF framework and integrated it into OAI’s core network framework. Then within that LMF framework, we develop the determine-location API [9, Section 6] to handle the positioning request following the 3gpp standard.
Figure 5: Determine Location Request [9, Figure 5.2.2.2.2-1]
The 3GPP outlines multiple identification methods [9, Section 6.1.6.2.2] for initiating the positioning of a target UE. In our implementation, we utilized the Subscription Permanent Identifier (SUPI) to identify the target UE. The SUPI functions as a permanent identifier for subscribers within the network, supporting various authentication, authorization, and identification tasks in 5G networks. Specifically, in our implementation, the SUPI is represented by the International Mobile Subscriber Identity (IMSI), a commonly used identifier in mobile networks. The key information elements included in the InputData to initiate the location request for our implementation are as follows:
@@
The LMF initiates the UL-TDOA positioning method by sending a TRP information request [7, Section 9.1.1.14] to obtain the necessary TRP information for UE positioning. The key information elements included in our implementation of the TRP information request, based on the list provided by 3GPP [7, Section 9.1.1.14], are as follows:
@@
All gNBs participating in the positioning procedure prepare a TRP information response and send it to the LMF, providing essential data such as configuration and location, which are crucial for accurate UE positioning. The key information elements included in our implementation of the TRP information response, based on the list provided by 3GPP [7, Section 9.1.1.15], are as follows:
@@
Message 4 Positioning Information Request: After receiving the TRP Information Responses from all gNBs involved in the positioning process, the LMF initiates the retrieval of UL-SRS configuration details for the target UE. It then sends a Positioning Information Request message to the serving gNB. In response, the serving gNB allocates the necessary resources for UL-SRS and configures the target device with the appropriate UL-SRS resource sets. The key information elements included in our implementation of the Positioning Information Request, as outlined by 3GPP [7, Section 9.1.1.10], are as follows:
@@
Message 5 Positioning Information Response: After processing the Positioning Information Request, the serving gNB prepares a Positioning Information Response that includes the UL SRS configuration for the target UE. In the current OAI RAN framework, the SRS configuration is predefined in the gNB settings and does not allow for dynamic adjustments for individual UEs. Consequently, when the serving gNB receives the Positioning Information Request, it disregards the specific configuration requested by the LMF and instead returns the existing UL SRS configuration that has already been applied to the target UE. The key information elements included in our implementation of the Positioning Information response, as outlined by 3GPP [7, Section 9.1.1.11], are as follows:
@@
Once the LMF is informed of the SRS configuration for the target UE, it initiates the activation of UE SRS transmission by sending an NRPPa Positioning Activation Request message to the serving gNB. The key information elements included in our implementation of the Positioning Activation Request, as outlined by 3GPP [7, Section 9.1.1.12], are as follows:
@@
Message 7 Positioning Activation Response: In the current OAI RAN framework, UE SRS transmission is preactivated, and real-time modifications to a UE’s SRS transmission are not supported. As a result, when the gNB receives the request, it generates a Positioning Activation Response and sends it back to the LMF, indicating that the target device is transmitting the SRS. The key information elements included in our implementation of the Positioning Activation Response, as outlined by 3GPP [7, Section 9.1.1.18], are as follows:
@@
Once the Positioning Activation Response is received by the LMF, it initiates the retrieval of ToA measurements from all the gNBs/TRPs involved in the positioning of the target UE. The LMF sends a Measurement Request to each participating gNB, which includes the SRS configuration previously received from the serving gNB in the Positioning Information Response for the target UE. The participating gNBs then use this configuration to perform the ToA measurements and send the results back to the LMF. The key information elements included in our implementation of the measurement Request, as outlined by 3GPP [7, Section 9.1.4.1], are as follows:
@@
All gNBs that have received the Measurement Request and are involved in the positioning procedure generate a Measurement Response and send it to the LMF, providing crucial ToA measurements necessary for accurate UE positioning. The precision and resolution of these ToA measurements are essential for achieving high positioning accuracy, as any errors in the ToA measurements can significantly affect the overall accuracy. To enhance measurement precision, we have integrated a state-of-the-art channel estimation and ToA estimation procedure into the gNB. A detailed overview of this integration is provided in Section III-B. The key information elements included in our implementation of the measurement response, based on the list provided by 3GPP [7, Section 9.1.4.2], are as follows:
@@
Once the LMF has calculated the position of the target UE, it initiates the process to provide a response to the External API that initiated the positioning request, as illustrated in Figure 5. This response includes a data structure of type LocationData [9, Section 6.1.6.2.3]. The key information elements included in our implementation of the positioning response (i.e., LocationData), based on the list provided by 3GPP [9, Section 6.1.6.2.3], are as follows:
@@ IV-A1 Preparing the Setup
The next step is to clone the OAI 5G RAN repository [15] and build the RFsimulator-based gNB and UE. Our work is part of the NRPPA_Procedures branch [^3] The main branch of the OAI 5G RAN repository is develop, and all new features are eventually merged into this branch. If the NRPPA_Procedures branch no longer exists, it indicates that the features have been merged into the develop branch of the OAI 5G RAN repository, and you should use the develop branch instead..
Once the NF images have been pulled and the RFsimulator-based gNB and UE have been successfully built, we are ready to deploy our RFsimulator-based setup using these components. Use the following commands to deploy the 5G core. Figure 7 shows the logs from a successful 5G core deployment.
@@
The logs of the AMF can be checked using the following command. Figure 8 shows the AMF logs prior to starting the gNB and UE.
@@
to the determine-location API of LMF, where the request contains a data structure of type InputData [9, Section 6.1.6.2.2] which contain the IMSI value of the UE. We can initiate the positioning resuest using the following command, where Using the following command, where the IP address of the LMF is $192.168.70.141$.
Now we are ready to start the positioning procedure for the UE. To initiate the positioning process for a specific UE, we send an HTTP POST request to the determine-location API of the LMF. The request contains a data structure of type InputData [9, Section 6.1.6.2.2], which includes the IMSI value of the UE. We can initiate the positioning request using the following command, where the IP address of the LMF is $192.168.70.141$.
@@
Figure 14: Example logs when LMF received TRP information request
V O-RAN based Localization Testbed at EURECOM
@@
V-A Deployment
In order to evaluate the localization features described above, we have deployed 3 additional O-RUs from Firecell outdoors (see Table I for details). An aerial photo of the EURECOM building is shown in Figure 15. The building has two symmetric wings and two different levels following the slope of the hill it is built on. The top level roof features two antenna sites, one on each wing, which are part of the greater EURECOM 5G testbed. Each site has multiple fibre connections that run directly into EURECOM’s server room which hosts the computing and switching infrastructure. We have deployed a Firecell RU on each of these sites and each RU drives 2 external antennas, which have been mounted on the railings of the roof and are connected to the RU through a 10m low loss cable (the cable introduces an attenuation of xdB, which has been included in the link budget). These antennas overlook the lower roof of the Eurecom building, which is three floors below. The northeast wing of the lower roof is accessible to people and is used as an experimentation area. The southwest wing of the lower roof is off-limits for people and features a third RU whose antennas are deployed on small tripods spaced 10m apart and at a height of 2m [to be checked]. Our deployment can be seen in Figures 15 and 17.
@@ V-B Logical Architecture
The CU and DU run on a server ”colibri” with a Intel(R) Xeon(R) Gold 6354 CPU @ 3.00GHz CPUs with 18 cores each. For the fronthaul network it uses an Intel X710 4x10Gbps NIC. Two of the four ports are used to connect to the backhaul network (via a CISCO C9364C-GX switch) and the two other ports are connected to the fronthaul network. The core network runs on the server ”alambix” in a docker environmet.
@@
V-C Ground truth measurements
@@
VI Results
This section describes the process of validating the overall performance of a deployed 5G network by testing its end-to-end functionality. The validation involves running the OAI 5G CN components, including the oai-lmf, alongside a gNB that integrates CU-DU functions, and a Firecell’s ORAN RU. A 5G phone, acting as a UE, is then registered with this network, and the success of this registration is verified by analyzing the logs generated by the oai-amf as below:
gNBs’ Information
Index Status Global Id gNB Name PLMN
1 Connected 0xE010 gNB-Geo5G 001,01
UEs’ Information
Index 5GMM State IMSI GUTI
1 Connected UE’s IMSI UE’s GUTI
Once we ensure the UE and gNB are connected we trigger the determine location API to start the positioning request.

### v2

@@ I Introduction
There are several examples where OAI is used as a basis for positioning. In [3], OAI is combined with a multi-channel SDR and a custom 2x2 patch antenna array for enhanced cell-ID (E-CID) positioning in indoor scenarios, by using RTT and AoA estimates based on the demodulation reference signal (DM-RS). The HOP-5G project [4, 5] sponsored the development of the PRS in OAI for DL-TDoA positioning for FR1 and FR2 and demonstrated the feasibility of this techniques. Based on this work, [6] built a testbed with three OAI base stations or gNBs and a OAI user equipment exploiting the PRS transmissions for DL-TDoA positioning. More recently, a novel framework is presented in [7] to estimate the RTT between a UE and a base station (gNB) with existing 5G NR transmissions.
@@ III-A UL-TDoA Positioning Procedure in OAI
The UL-TDoA positioning procedure begins with an external API generating a location information request for a target UE for the LMF. As shown in Figure  5 to initiate the location procedure, the external API (which can be an advanced API or a simple one-line command) sends the HTTP post request to determine-location API of LMF, where the request contains a data structure of type InputData [10, Section 6.1.6.2.2]. To enable the exchange of this message, we first developed the 3gpp-compliant LMF framework and integrated it into OAI’s core network framework. Then within that LMF framework, we develop the determine-location API [10, Section 6] to handle the positioning request following the 3gpp standard.
Figure 5: Determine Location Request [10, Figure 5.2.2.2.2-1]
The 3GPP outlines multiple identification methods [10, Section 6.1.6.2.2] for initiating the positioning of a target UE. In our implementation, we utilized the Subscription Permanent Identifier (SUPI) to identify the target UE. The SUPI functions as a permanent identifier for subscribers within the network, supporting various authentication, authorization, and identification tasks in 5G networks. Specifically, in our implementation, the SUPI is represented by the International Mobile Subscriber Identity (IMSI), a commonly used identifier in mobile networks. The key information elements included in the InputData to initiate the location request for our implementation are as follows:
@@
The LMF initiates the UL-TDOA positioning method by sending a TRP information request [8, Section 9.1.1.14] to obtain the necessary TRP information for UE positioning. The key information elements included in our implementation of the TRP information request, based on the list provided by 3GPP [8, Section 9.1.1.14], are as follows:
@@
All gNBs participating in the positioning procedure prepare a TRP information response and send it to the LMF, providing essential data such as configuration and location, which are crucial for accurate UE positioning. The key information elements included in our implementation of the TRP information response, based on the list provided by 3GPP [8, Section 9.1.1.15], are as follows:
@@
Message 4 Positioning Information Request: After receiving the TRP Information Responses from all gNBs involved in the positioning process, the LMF initiates the retrieval of UL-SRS configuration details for the target UE. It then sends a Positioning Information Request message to the serving gNB. In response, the serving gNB allocates the necessary resources for UL-SRS and configures the target device with the appropriate UL-SRS resource sets. The key information elements included in our implementation of the Positioning Information Request, as outlined by 3GPP [8, Section 9.1.1.10], are as follows:
@@
Message 5 Positioning Information Response: After processing the Positioning Information Request, the serving gNB prepares a Positioning Information Response that includes the UL SRS configuration for the target UE. In the current OAI RAN framework, the SRS configuration is predefined in the gNB settings and does not allow for dynamic adjustments for individual UEs. Consequently, when the serving gNB receives the Positioning Information Request, it disregards the specific configuration requested by the LMF and instead returns the existing UL SRS configuration that has already been applied to the target UE. The key information elements included in our implementation of the Positioning Information response, as outlined by 3GPP [8, Section 9.1.1.11], are as follows:
@@
Once the LMF is informed of the SRS configuration for the target UE, it initiates the activation of UE SRS transmission by sending an NRPPa Positioning Activation Request message to the serving gNB. The key information elements included in our implementation of the Positioning Activation Request, as outlined by 3GPP [8, Section 9.1.1.12], are as follows:
@@
Message 7 Positioning Activation Response: In the current OAI RAN framework, UE SRS transmission is preactivated, and real-time modifications to a UE’s SRS transmission are not supported. As a result, when the gNB receives the request, it generates a Positioning Activation Response and sends it back to the LMF, indicating that the target device is transmitting the SRS. The key information elements included in our implementation of the Positioning Activation Response, as outlined by 3GPP [8, Section 9.1.1.18], are as follows:
@@
Once the Positioning Activation Response is received by the LMF, it initiates the retrieval of ToA measurements from all the gNBs/TRPs involved in the positioning of the target UE. The LMF sends a Measurement Request to each participating gNB, which includes the SRS configuration previously received from the serving gNB in the Positioning Information Response for the target UE. The participating gNBs then use this configuration to perform the ToA measurements and send the results back to the LMF. The key information elements included in our implementation of the measurement Request, as outlined by 3GPP [8, Section 9.1.4.1], are as follows:
@@
All gNBs that have received the Measurement Request and are involved in the positioning procedure generate a Measurement Response and send it to the LMF, providing crucial ToA measurements necessary for accurate UE positioning. The precision and resolution of these ToA measurements are essential for achieving high positioning accuracy, as any errors in the ToA measurements can significantly affect the overall accuracy. To enhance measurement precision, we have integrated a state-of-the-art channel estimation and ToA estimation procedure into the gNB. A detailed overview of this integration is provided in Section III-B. The key information elements included in our implementation of the measurement response, based on the list provided by 3GPP [8, Section 9.1.4.2], are as follows:
@@
Once the LMF has calculated the position of the target UE, it initiates the process to provide a response to the External API that initiated the positioning request, as illustrated in Figure 5. This response includes a data structure of type LocationData [10, Section 6.1.6.2.3]. The key information elements included in our implementation of the positioning response (i.e., LocationData), based on the list provided by 3GPP [10, Section 6.1.6.2.3], are as follows:
@@ IV-A1 Preparing the Setup
#Pull OAI Core Network Functionsdocker pull oaisoftwarealliance/ims:latestdocker pull oaisoftwarealliance/oai-amf:developdocker pull oaisoftwarealliance/oai-nrf:developdocker pull oaisoftwarealliance/oai-smf:developdocker pull oaisoftwarealliance/oai-udr:developdocker pull oaisoftwarealliance/oai-upf:developdocker pull oaisoftwarealliance/oai-udm:developdocker pull oaisoftwarealliance/oai-ausf:developdocker pull oaisoftwarealliance/oai-lmf:developdocker pull oaisoftwarealliance/trf-gen-cn5g:latest The next step is to clone the OAI 5G RAN repository [16] and build the RFsimulator-based gNB and UE. Our work is part of the NRPPA_Procedures branch [^3] The main branch of the OAI 5G RAN repository is develop, and all new features are eventually merged into this branch. If the NRPPA_Procedures branch no longer exists, it indicates that the features have been merged into the develop branch of the OAI 5G RAN repository, and you should use the develop branch instead..
 git clone https://gitlab.eurecom.fr/oai/openairinterface5g.git cd openairinterface5g/cmake_targets git checkout NRPPA_Procedures # install dependencies ./build_oai -I # compile gNB and nrUE ./build_oai --gNB --nrUE -w SIMU
Once the NF images have been pulled and the RFsimulator-based gNB and UE have been successfully built, we are ready to deploy our RFsimulator-based setup using these components. Use the following commands to deploy the 5G core.
cd openairinterface5g/doc/tutorial_resources/oai-cn5gdocker compose -f docker-compose.yaml up -d# verify the deploymentdocker ps -a
Figure 7 shows the logs from a successful 5G core deployment.
@@
The logs of the AMF can be checked using the following command.
 docker logs oai-amf -f
Figure 8 shows the AMF logs prior to starting the gNB and UE.
@@
#open a terminal to run the gNBcd openairinterface5g/cmake_targets/ran_build/buildsudo RFSIMULATOR=server ./nr-softmodem -O ../../../targets/PROJECTS/GENERIC-NR-5GC/CONF/ gnb.sa.band78.fr1.106PRB.usrpb210.conf --gNBs.[0].min_rxtxtime 6 --rfsim --sa#open another terminal to run the UEcd openairinterface5g/cmake_targets/ran_build/buildsudo RFSIMULATOR=127.0.0.1 ./nr-uesoftmodem -r 106 --numerology 1 --band 78 -C 3619200000 --sa --uicc0.imsi 001010000000001 --rfsim
@@
Now we are ready to start the positioning procedure for the UE. To initiate the positioning process for a specific UE, we send an HTTP POST request to the determine-location API of the LMF. The request contains a data structure of type InputData [10, Section 6.1.6.2.2], which includes the IMSI value of the UE. We can initiate the positioning request using the following command, where the IP address of the LMF is $192.168.70.141$.
 curl --http2-prior-knowledge -H "Content-Type: application/json" -d "@InputData.json" -X POST http://192.168.70.141:8080/nlmf-loc/v1/determine-location
@@
Figure 14: Example logs when LMF received ToA information
IV-B O-RAN based Localization Testbed at EURECOM
@@
IV-B1 Deployment
In order to evaluate the localization features described above, we have deployed 3 additional O-RUs from Firecell outdoors (see Table I for details). An aerial photo of the EURECOM building is shown in Figure 15. The building has two symmetric wings and two different levels following the slope of the hill it is built on. The top level roof features two antenna sites, one on each wing, which are part of the greater EURECOM 5G testbed. Each site has multiple fibre connections that run directly into EURECOM’s server room which hosts the computing and switching infrastructure. We have deployed a Firecell RU on each of these sites and each RU drives 2 external antennas, which have been mounted on the railings of the roof and are connected to the RU through a 10m low loss cable (the cable introduces an attenuation of 7 dB, which has been included in the link budget). These antennas overlook the lower roof of the Eurecom building, which is three floors below. The northeast wing of the lower roof is accessible to people and is used as an experimentation area. The southwest wing of the lower roof is off-limits for people and features a third RU whose antennas are deployed on small tripods spaced 10m apart and at a height of 2m. Our deployment can be seen in Figures 15 and 17.
@@ IV-B2 Logical Architecture
The CU and DU run on a server ”colibri” with a Intel(R) Xeon(R) Gold 6354 CPU @ 3.00GHz CPUs with 18 cores each. For the fronthaul network it uses an Intel X710 4x10Gbps NIC. Two of the four ports are used to connect to the backhaul network (via a CISCO C9364C-GX switch) and the two other ports are connected to the fronthaul network. The core network runs on the server ”alambix” in a docker environment.
@@
IV-B3 Ground truth measurements
@@
IV-C Results
This section describes the process of validating the overall performance of a deployed 5G network by testing its end-to-end functionality. The validation involves running the OAI 5G CN components, including the oai-lmf, alongside a gNB that integrates CU-DU functions, and a Firecell’s ORAN RU. A 5G phone, acting as a UE, is then registered with this network, and the success of this registration is verified by analyzing the logs generated by the oai-amf as shown in Figure 10. Once we ensure the UE and gNB are connected we trigger the determine location API to start the positioning request.